OpenStudy (anonymous):

how do you plus in x=3cos^3(theta) and y=3sin^3(theta) into a graphing calc. I have trouble with raising cosine and sine to the third

OpenStudy (ryan123345):

Hey there! I see that you are new to OpenStudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! If you have any additional questions involving the website feel free to ask the people with the purples A's next to their name. They are the ambassadors and will answer all your questions regarding the site. And also read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Have a great day and I hope you have fun!

OpenStudy (anonymous):

Does your calculator allow parametrized curves?

OpenStudy (anonymous):

\[ \sin^3(\theta) = [\sin(\theta)]^3 \]